# Rotating black holes in de Rham-Gabadadze-Tolley massive gravity: Analytic calculation procedure

### v2

@@ V The ansätz of Stückelberg fields $\phi^{a}$
where 00 is a $2\times 2$ zero matrix and
@@
$M^{2}_{11}$ $=\frac{1}{2}\bigg{(}[M]^{2}-2\text{det}~{}M-\sqrt{[M]^{4}-4[M]^{2}\text{det}~{}M-4M^{2}_{12}M^{2}_{21}}\bigg{)},$ (65)
$M^{2}_{22}$ $=\frac{1}{2}\bigg{(}[M]^{2}-2\text{det}~{}M+\sqrt{[M]^{4}-4[M]^{2}\text{det}~{}M-4M^{2}_{12}M^{2}_{21}}\bigg{)}.$ (66)
@@
$s_{1}$ $=\sqrt{[\sqrt{\Xi_{1}}]^{4}-4[\sqrt{\Xi_{1}}]^{2}\text{det}~{}\sqrt{\Xi_{1}}+4e^{-2\nu}(f_{01}+qf_{11})(e^{-2\psi}f_{10}-e^{-2\nu}q(f_{00}+qf_{01}))},$ (73)
@@
$T^{(\mathcal{K})}{}^{0}{}_{0}$ $=\frac{1}{2[\sqrt{\Xi_{1}}]}\bigg{(}[\sqrt{\Xi_{1}}](-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}]))$
$+([\sqrt{\Xi_{1}}]^{2}+s_{1}+2e^{-2\nu}(f_{01}+qf_{11})q)(1+c_{3}(\text{det}~{}\sqrt{\Xi_{2}}-1)+c_{4}(2-[\sqrt{\Xi_{2}}]))\bigg{)},$ (77)
$T^{(\mathcal{K})}{}^{1}{}_{1}$ $=\frac{1}{2[\sqrt{\Xi_{1}}]}\bigg{(}[\sqrt{\Xi_{1}}](-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}]))$
$+([\sqrt{\Xi_{1}}]^{2}-s_{1}-2e^{-2\nu}(f_{01}+qf_{11})q)(1+c_{3}(\text{det}~{}\sqrt{\Xi_{2}}-1)+c_{4}(2-[\sqrt{\Xi_{2}}]))\bigg{)},$ (78)
$T^{(\mathcal{K})}{}^{2}{}_{2}$ $=\frac{1}{2[\sqrt{\Xi_{2}}]}\bigg{(}[\sqrt{\Xi_{2}}](-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}]))$
$+([\sqrt{\Xi_{2}}]^{2}+s_{2})(1+c_{3}(\text{det}~{}\sqrt{\Xi_{1}}-1)+c_{4}(2-[\sqrt{\Xi_{1}}]))\bigg{)},$ (79)
$T^{(\mathcal{K})}{}^{3}{}_{3}$ $=\frac{1}{2[\sqrt{\Xi_{2}}]}\bigg{(}[\sqrt{\Xi_{2}}](-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}]))$
$+([\sqrt{\Xi_{2}}]^{2}-s_{2})(1+c_{3}(\text{det}~{}\sqrt{\Xi_{1}}-1)+c_{4}(2-[\sqrt{\Xi_{1}}]))\bigg{)},$ (80)
$T^{(\mathcal{K})}{}^{0}{}_{1}$ $=\frac{(f_{01}+qf_{11})e^{-\nu-\psi}}{[\sqrt{\Xi_{1}}]}(1+c_{3}(\text{det}~{}\sqrt{\Xi_{2}}-1)+c_{4}(2-[\sqrt{\Xi_{2}}])),$ (81)
$T^{(\mathcal{K})}{}^{2}{}_{3}$ $=-\frac{f_{23}e^{-\mu_{2}-\mu_{3}}}{[\sqrt{\Xi_{2}}]}(1+c_{3}(\text{det}~{}\sqrt{\Xi_{1}}-1)+c_{4}(2-[\sqrt{\Xi_{1}}])).$ (82)
V.1 the cosmological constant term $\Lambda$
@@
$-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}])$ $=2\Lambda,$ (85)
$-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)+c_{4}(-3-2\text{det}~{}\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}])$ $=2\Lambda,$ (86)
@@
$\text{det}~{}\sqrt{\Xi_{1}}$ $=\text{det}~{}\sqrt{\Xi_{2}}=-\frac{3c_{3}-2c_{3}^{2}-2c_{4}+5c_{3}c_{4}-4c_{4}^{2}+2c_{3}\Lambda}{2(-c_{3}+c_{3}^{2}-2c_{3}c_{4}+c_{4}^{2})},$ (87)
$[\sqrt{\Xi_{1}}]$ $=[\sqrt{\Xi_{2}}]=-\frac{-2+4c_{3}-2c_{3}^{2}-5c_{4}+6c_{3}c_{4}-5c_{4}^{2}+2c_{4}\Lambda}{2(-c_{3}+c_{3}^{2}-2c_{3}c_{4}+c_{4}^{2})}.$ (88)
@@
$T^{(\mathcal{K})}_{S}{}^{0}{}_{0}$ $=-T^{(\mathcal{K})}_{S}{}^{1}{}_{1}=-e^{-2\psi-2\mu_{2}}(A^{\prime 2}_{,2}+B^{\prime 2}_{,2})-e^{-2\psi-2\mu_{3}}(A^{\prime 2}_{,3}+B^{\prime 2}_{,3}),$ (90)
$T^{(\mathcal{K})}_{S}{}^{2}{}_{2}$ $=-T^{(\mathcal{K})}_{S}{}^{3}{}_{3}=e^{-2\psi-2\mu_{2}}(A^{\prime 2}_{,2}+B^{\prime 2}_{,2})-e^{-2\psi-2\mu_{3}}(A^{\prime 2}_{,3}+B^{\prime 2}_{,3}),$ (91)
@@
VI Gauge freedom and the Field equations
Analytical calculation of the field equations (10) is not an easy task. The original mathematical process is detailed in Refs. [27, 28]. A more streamlined calculation process is provided by Carter [29] or Chandrasekhar [30]. In this paper, we employ the methodology primarily adopted from Chandrasekhar. As a key strategy of Chandrasekhar, we require a highly symmetric metirc form [26]
@@ VI Gauge freedom and the Field equations
$G^{2}{}_{2}+G^{3}{}_{3}+m^{2}(T^{(\mathcal{K})}{}^{2}{}_{2}+T^{(\mathcal{K})}{}^{3}{}_{3})$ $=8\pi(T^{(m)}{}^{2}{}_{2}+T^{(m)}{}^{3}{}_{3}),$ (100)
$G^{2}{}_{2}-G^{3}{}_{3}+m^{2}(T^{(\mathcal{K})}{}^{2}{}_{2}-T^{(\mathcal{K})}{}^{3}{}_{3})$ $=8\pi(T^{(m)}{}^{2}{}_{2}-T^{(m)}{}^{3}{}_{3}),$ (101)
$G^{1}{}_{1}+G^{0}{}_{0}+m^{2}(T^{(\mathcal{K})}{}^{1}{}_{1}+T^{(\mathcal{K})}{}^{0}{}_{0})$ $=8\pi(T^{(m)}{}^{1}{}_{1}+T^{(m)}{}^{0}{}_{0}),$ (102)
$G^{1}{}_{1}-G^{0}{}_{0}+m^{2}(T^{(\mathcal{K})}{}^{1}{}_{1}-T^{(\mathcal{K})}{}^{0}{}_{0})$ $=8\pi(T^{(m)}{}^{1}{}_{1}-T^{(m)}{}^{0}{}_{0}),$ (103)
@@
$\Delta_{r,rr}-\Delta_{\mu,\mu\mu}-\frac{1}{4}\frac{\Delta_{r,r}^{2}}{\Delta_{r}}+\frac{1}{4}\frac{\Delta_{\mu,\mu}^{2}}{\Delta_{\mu}}-\Delta_{r,r}(\ln\rho^{2})_{,r}+\Delta_{\mu,\mu}(\ln\rho^{2})_{,\mu}$
$+\frac{\Delta_{r}}{\chi^{2}}(\chi^{2}_{,r}-q^{2}_{,r})-\frac{\Delta_{\mu}}{\chi^{2}}(\chi^{2}_{,\mu}-q^{2}_{,\mu})+2\chi\bigg{(}\sqrt{\frac{\Delta_{\mu}}{\Delta_{r}}}(A_{,\mu}^{2}-m^{2}A^{\prime}{}_{,\mu}^{2}+B_{,\mu}^{2}-m^{2}B^{\prime}{}_{,\mu}^{2})$
@@
$\Delta_{r,rr}+\Delta_{\mu,\mu\mu}-\frac{1}{4}\frac{\Delta_{r,r}^{2}}{\Delta_{r}}-\frac{1}{4}\frac{\Delta_{\mu,\mu}^{2}}{\Delta_{\mu}}+2\Delta_{r}(\ln\rho^{2})_{,rr}+2\Delta_{\mu}(\ln\rho^{2})_{,\mu\mu}$
$+\Delta_{r,r}(\ln\rho^{2})_{,r}+\Delta_{\mu,\mu}(\ln\rho^{2})_{,\mu}+\frac{\Delta_{r}}{\chi^{2}}(\chi^{2}_{,r}-q^{2}_{,r})+\frac{\Delta_{\mu}}{\chi^{2}}(\chi^{2}_{,\mu}-q^{2}_{,\mu})+m^{2}\Lambda\rho^{2}$ $=0,$ (107)
@@
$+2\chi\bigg{(}\sqrt{\frac{\Delta_{\mu}}{\Delta_{r}}}(A_{,\mu}^{2}-m^{2}A^{\prime}{}_{,\mu}^{2}+B_{,\mu}^{2}-m^{2}B^{\prime}{}_{,\mu}^{2})$
@@
$\left(\frac{\Delta_{r}}{\chi^{2}}q_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\chi^{2}}q_{,\mu}\right)_{,\mu}-4\bigg{(}B_{,r}A_{,\mu}-B_{,\mu}A_{,r}+m^{2}B^{\prime}_{,\mu}A^{\prime}_{,r}-m^{2}B^{\prime}_{,r}A^{\prime}_{,\mu}\bigg{)}$ $=0.$ (109)
@@
VII the Ernst equations
Similar to Ref. [26], we solve Eqs. (105) and (108) - (113) simultaneously. We rewrite the pairs $A,B$ into a single complex potential
@@
where * denotes the conjugate of complex numbers and ||||| | denotes the modular length of a complex number. Then, we can reexpressed Eqs. (108) and (109) as
@@
$\left(\frac{\Delta_{r}}{\chi^{2}}q_{,r}-2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})\right)_{,r}+\left(\frac{\Delta_{\mu}}{\chi^{2}}q_{,\mu}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r})\right)_{,\mu}$ $=0,$ (120)
@@
$\left(\frac{\chi^{2}}{\Delta_{\mu}}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))\right)_{,r}+\left(\frac{\chi^{2}}{\Delta_{r}}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))\right)_{,\mu}=0.$ (123)
@@
$\Psi\{[\Delta_{r}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}]_{,r}+[\Delta_{\mu}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}]_{,\mu}\}=\frac{1}{2}\Psi^{2}(\Delta_{r,rr}+\Delta_{\mu,\mu\mu})$
$+\Delta_{r}\{\Psi_{,r}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}-\Phi_{,r}^{2}-2\Phi_{,r}\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Psi_{,\mu}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}-\Phi_{,\mu}^{2}-2\Phi_{,\mu}\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})\},$ (125)
$\Psi[(\Delta_{r}\Phi_{,r})_{,r}+(\Delta_{\mu}\Phi_{,\mu})_{,\mu}]=$
$\Delta_{r}\{\Phi_{,r}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}\text{Im}(HH^{*}_{,r}-VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Phi_{,\mu}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu})\},$ (126)
@@
$\Psi[(\Delta_{r}H_{,r})_{,r}+(\Delta_{\mu}H_{,\mu})_{,\mu}]$ $=\Delta_{r}H_{,r}[\Psi_{,r}+i(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))]$
$+\Delta_{\mu}H_{,\mu}[\Psi_{,\mu}+i(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))].$ (127)
@@
$\Psi[(\Delta_{r}V_{,r})_{,r}+(\Delta_{\mu}V_{,\mu})_{,\mu}]$ $=\Delta_{r}V_{,r}[\Psi_{,r}+i(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))]$
$+\Delta_{\mu}V_{,\mu}[\Psi_{,\mu}+i(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))].$ (128)
@@
$\Psi[(\Delta_{r}Z_{,r})_{,r}+(\Delta_{\mu}Z_{,\mu})_{,\mu}]$ $=\frac{1}{2}\Psi^{2}(\Delta_{r,rr}+\Delta_{\mu,\mu\mu})+\Delta_{r}Z_{,r}[Z_{,r}-2(H^{*}H_{,r}-m^{2}V^{*}V_{,r})]$
@@
$+2\tilde{\chi}\bigg{(}\sqrt{\frac{\Delta_{\mu}}{\Delta_{r}}}(\tilde{A}_{,\mu}^{2}-m^{2}\tilde{A}^{\prime}{}_{,\mu}^{2}+\tilde{B}_{,\mu}^{2}-m^{2}\tilde{B}^{\prime}{}_{,\mu}^{2})$
@@
$\left(\frac{\Delta_{r}}{\tilde{\chi}^{2}}\tilde{q}_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\tilde{\chi}^{2}}\tilde{q}_{,\mu}\right)_{,\mu}-4\bigg{(}\tilde{B}_{,r}\tilde{A}_{,\mu}-\tilde{B}_{,\mu}\tilde{A}_{,r}+m^{2}\tilde{B}^{\prime}_{,\mu}\tilde{A}^{\prime}_{,r}-m^{2}\tilde{B}^{\prime}_{,r}\tilde{A}^{\prime}_{,\mu}\bigg{)}$ $=0,$
@@
IX the homogeneous solution
Similar to the Ref. [26], we first solve the Eqs. (105) and (131) - (134) in the homogeneous case, which is $\Lambda=0$. We denote the homogeneous solution with a corner ”K”. When difine
@@
$\{[\frac{1}{2}$ $-(|Q|^{2}-m^{2}|S|^{2})](Z^{K}+(Z^{K})^{*})-(|Q|^{2}-m^{2}|S|^{2})(|Z^{K}|^{2}+1)\}[(\Delta_{r}^{K}(Z^{K})_{,r})_{,r}+(\Delta_{\mu}^{K}(Z^{K})_{,\mu})_{,\mu}]$
=$=\{1-2(|Q|^{2}-m^{2}|S|^{2})((Z^{K})^{*}+1)\}[\Delta_{r}^{K}(Z^{K})_{,r}^{2}+\Delta_{\mu}^{K}(Z^{K})_{,\mu}^{2}],$ (150)

### v3

@@ 5 The ansätz of Stückelberg fields $\phi^{a}$
where 0 is a $2\times 2$ zero matrix and
@@
$M^{2}_{11}$ $=\frac{1}{2}\bigg{(}[M]^{2}-2\text{det}~M$
$-\sqrt{[M]^{4}-4[M]^{2}\text{det}~M-4M^{2}_{12}M^{2}_{21}}\bigg{)},$ (65)
$M^{2}_{22}$ $=\frac{1}{2}\bigg{(}[M]^{2}-2\text{det}~M$
$+\sqrt{[M]^{4}-4[M]^{2}\text{det}~M-4M^{2}_{12}M^{2}_{21}}\bigg{)}.$ (66)
@@
$s_{1}$ $=\big{(}[\sqrt{\Xi_{1}}]^{4}-4[\sqrt{\Xi_{1}}]^{2}\text{det}~\sqrt{\Xi_{1}}+4e^{-2\nu}(f_{01}$
$+qf_{11})(e^{-2\psi}f_{10}-e^{-2\nu}q(f_{00}+qf_{01}))\big{)}^{1/2},$ (73)
@@
$T^{(\mathcal{K})}{}^{0}{}_{0}$ $=\frac{1}{2[\sqrt{\Xi_{1}}]}\bigg{(}[\sqrt{\Xi_{1}}](-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}]))+([\sqrt{\Xi_{1}}]^{2}+s_{1}$
$+2e^{-2\nu}(f_{01}+qf_{11})q)(1+c_{3}(\text{det}~\sqrt{\Xi_{2}}-1)$
$+c_{4}(2-[\sqrt{\Xi_{2}}]))\bigg{)},$ (77)
$T^{(\mathcal{K})}{}^{1}{}_{1}$ $=\frac{1}{2[\sqrt{\Xi_{1}}]}\bigg{(}[\sqrt{\Xi_{1}}](-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}]))+([\sqrt{\Xi_{1}}]^{2}-s_{1}$
$-2e^{-2\nu}(f_{01}+qf_{11})q)(1+c_{3}(\text{det}~\sqrt{\Xi_{2}}-1)$
$+c_{4}(2-[\sqrt{\Xi_{2}}]))\bigg{)},$ (78)
$T^{(\mathcal{K})}{}^{2}{}_{2}$ $=\frac{1}{2[\sqrt{\Xi_{2}}]}\bigg{(}[\sqrt{\Xi_{2}}](-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}]))+([\sqrt{\Xi_{2}}]^{2}$
$+s_{2})(1+c_{3}(\text{det}~\sqrt{\Xi_{1}}-1)+c_{4}(2-[\sqrt{\Xi_{1}}]))\bigg{)},$ (79)
$T^{(\mathcal{K})}{}^{3}{}_{3}$ $=\frac{1}{2[\sqrt{\Xi_{2}}]}\bigg{(}[\sqrt{\Xi_{2}}](-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}]))+([\sqrt{\Xi_{2}}]^{2}$
$-s_{2})(1+c_{3}(\text{det}~\sqrt{\Xi_{1}}-1)+c_{4}(2-[\sqrt{\Xi_{1}}]))\bigg{)},$ (80)
$T^{(\mathcal{K})}{}^{0}{}_{1}$ $=\frac{(f_{01}+qf_{11})e^{-\nu-\psi}}{[\sqrt{\Xi_{1}}]}(1+c_{3}(\text{det}~\sqrt{\Xi_{2}}-1)$
$+c_{4}(2-[\sqrt{\Xi_{2}}])),$ (81)
$T^{(\mathcal{K})}{}^{2}{}_{3}$ $=-\frac{f_{23}e^{-\mu_{2}-\mu_{3}}}{[\sqrt{\Xi_{2}}]}(1+c_{3}(\text{det}~\sqrt{\Xi_{1}}-1)$
$+c_{4}(2-[\sqrt{\Xi_{1}}])).$ (82)
5.1 the cosmological constant term $\Lambda$
@@
$-3+2[\sqrt{\Xi_{2}}]-2c_{3}([\sqrt{\Xi_{2}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{2}}+4[\sqrt{\Xi_{2}}])$ $=2\Lambda,$ (85)
$-3+2[\sqrt{\Xi_{1}}]-2c_{3}([\sqrt{\Xi_{1}}]-1)$
$+c_{4}(-3-2\text{det}~\sqrt{\Xi_{1}}+4[\sqrt{\Xi_{1}}])$ $=2\Lambda,$ (86)
@@
$\text{det}~\sqrt{\Xi_{1}}$ $=\text{det}~\sqrt{\Xi_{2}}=$
$-\frac{3c_{3}-2c_{3}^{2}-2c_{4}+5c_{3}c_{4}-4c_{4}^{2}+2c_{3}\Lambda}{2(-c_{3}+c_{3}^{2}-2c_{3}c_{4}+c_{4}^{2})},$ (87)
$[\sqrt{\Xi_{1}}]$ $=[\sqrt{\Xi_{2}}]=$
$-\frac{-2+4c_{3}-2c_{3}^{2}-5c_{4}+6c_{3}c_{4}-5c_{4}^{2}+2c_{4}\Lambda}{2(-c_{3}+c_{3}^{2}-2c_{3}c_{4}+c_{4}^{2})}.$ (88)
@@
$T^{(\mathcal{K})}_{S}{}^{0}{}_{0}$ $=-T^{(\mathcal{K})}_{S}{}^{1}{}_{1}=-e^{-2\psi-2\mu_{2}}(A^{\prime 2}_{,2}+B^{\prime 2}_{,2})$
$-e^{-2\psi-2\mu_{3}}(A^{\prime 2}_{,3}+B^{\prime 2}_{,3}),$ (90)
$T^{(\mathcal{K})}_{S}{}^{2}{}_{2}$ $=-T^{(\mathcal{K})}_{S}{}^{3}{}_{3}=e^{-2\psi-2\mu_{2}}(A^{\prime 2}_{,2}+B^{\prime 2}_{,2})$
$-e^{-2\psi-2\mu_{3}}(A^{\prime 2}_{,3}+B^{\prime 2}_{,3}),$ (91)
@@
6 Gauge freedom and the Field equations
Analytical calculation of the field equations (10) is not an easy task. The original mathematical process is detailed in Refs. Newman1965 ; Debney1969 . A more streamlined calculation process is provided by Carter Carter2009 or Chandrasekhar Chandrasekhar:1983 . In this paper, we employ the methodology primarily adopted from Chandrasekhar. As a key strategy of Chandrasekhar, we require a highly symmetric metirc form Li2024-2
@@ 6 Gauge freedom and the Field equations
[…]
@@
$\Delta_{r,rr}-\Delta_{\mu,\mu\mu}-\frac{1}{4}\frac{\Delta_{r,r}^{2}}{\Delta_{r}}+\frac{1}{4}\frac{\Delta_{\mu,\mu}^{2}}{\Delta_{\mu}}-\Delta_{r,r}(\ln\rho^{2})_{,r}$
$+\Delta_{\mu,\mu}(\ln\rho^{2})_{,\mu}+\frac{\Delta_{r}}{\chi^{2}}(\chi^{2}_{,r}-q^{2}_{,r})-\frac{\Delta_{\mu}}{\chi^{2}}(\chi^{2}_{,\mu}-q^{2}_{,\mu})$
[…]
@@
$\Delta_{r,rr}+\Delta_{\mu,\mu\mu}-\frac{1}{4}\frac{\Delta_{r,r}^{2}}{\Delta_{r}}-\frac{1}{4}\frac{\Delta_{\mu,\mu}^{2}}{\Delta_{\mu}}+2\Delta_{r}(\ln\rho^{2})_{,rr}$
$+2\Delta_{\mu}(\ln\rho^{2})_{,\mu\mu}+\Delta_{r,r}(\ln\rho^{2})_{,r}+\Delta_{\mu,\mu}(\ln\rho^{2})_{,\mu}$
$+\frac{\Delta_{r}}{\chi^{2}}(\chi^{2}_{,r}-q^{2}_{,r})+\frac{\Delta_{\mu}}{\chi^{2}}(\chi^{2}_{,\mu}-q^{2}_{,\mu})+m^{2}\Lambda\rho^{2}=0$ , (107)
@@
[…]
@@
$\left(\frac{\Delta_{r}}{\chi^{2}}q_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\chi^{2}}q_{,\mu}\right)_{,\mu}-4\bigg{(}B_{,r}A_{,\mu}-B_{,\mu}A_{,r}$
$+m^{2}B^{\prime}_{,\mu}A^{\prime}_{,r}-m^{2}B^{\prime}_{,r}A^{\prime}_{,\mu}\bigg{)}=0$ . (109)
@@
7 the Ernst equations
Similar to Ref. Li2024-2 , we solve Eqs. (105) and (108) - (113) simultaneously. We rewrite the pairs $A,B$ into a single complex potential
@@
where * denotes the conjugate of complex numbers and $||$ denotes the modular length of a complex number. Then, we can reexpressed Eqs. (108) and (109) as
@@
$\left(\frac{\Delta_{r}}{\chi^{2}}q_{,r}-2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})\right)_{,r}$
$+\left(\frac{\Delta_{\mu}}{\chi^{2}}q_{,\mu}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r})\right)_{,\mu}=0,$ (120)
@@
$\left(\frac{\chi^{2}}{\Delta_{\mu}}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))\right)_{,r}$
$+\left(\frac{\chi^{2}}{\Delta_{r}}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))\right)_{,\mu}=0.$ (123)
@@
$\Psi\{[\Delta_{r}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}]_{,r}+[\Delta_{\mu}(\Psi+|H|^{2}$
$-m^{2}|V|^{2})_{,\mu}]_{,\mu}\}=\frac{1}{2}\Psi^{2}(\Delta_{r,rr}+\Delta_{\mu,\mu\mu})+\Delta_{r}\{\Psi_{,r}(\Psi+|H|^{2}$
$-m^{2}|V|^{2})_{,r}-\Phi_{,r}^{2}-2\Phi_{,r}\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Psi_{,\mu}(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}$
$-\Phi_{,\mu}^{2}-2\Phi_{,\mu}\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})\},$ (125)
$\Psi[(\Delta_{r}\Phi_{,r})_{,r}+(\Delta_{\mu}\Phi_{,\mu})_{,\mu}]=\Delta_{r}\{\Phi_{,r}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}$
$+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}\text{Im}(HH^{*}_{,r}-VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Phi_{,\mu}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}$
$+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu})\},$ (126)
@@
$\Psi[(\Delta_{r}H_{,r})_{,r}+(\Delta_{\mu}H_{,\mu})_{,\mu}]=\Delta_{r}H_{,r}[\Psi_{,r}+i(\Phi_{,r}$
$+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))]+\Delta_{\mu}H_{,\mu}[\Psi_{,\mu}+i(\Phi_{,\mu}$
$+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))].$ (127)
@@
$\Psi[(\Delta_{r}V_{,r})_{,r}+(\Delta_{\mu}V_{,\mu})_{,\mu}]=\Delta_{r}V_{,r}[\Psi_{,r}+i(\Phi_{,r}$
$+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))]+\Delta_{\mu}V_{,\mu}[\Psi_{,\mu}$
$+i(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))].$ (128)
@@
$\Psi[(\Delta_{r}Z_{,r})_{,r}+(\Delta_{\mu}Z_{,\mu})_{,\mu}]=\frac{1}{2}\Psi^{2}(\Delta_{r,rr}+\Delta_{\mu,\mu\mu})$
$+\Delta_{r}Z_{,r}[Z_{,r}-2(H^{*}H_{,r}-m^{2}V^{*}V_{,r})]$
@@
[…]
@@
$\left(\frac{\Delta_{r}}{\tilde{\chi}^{2}}\tilde{q}_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\tilde{\chi}^{2}}\tilde{q}_{,\mu}\right)_{,\mu}-4\bigg{(}\tilde{B}_{,r}\tilde{A}_{,\mu}$
$-\tilde{B}_{,\mu}\tilde{A}_{,r}+m^{2}\tilde{B}^{\prime}_{,\mu}\tilde{A}^{\prime}_{,r}-m^{2}\tilde{B}^{\prime}_{,r}\tilde{A}^{\prime}_{,\mu}\bigg{)}=0,$
@@
9 the homogeneous solution
Similar to the Ref. Li2024-2 , we first solve the Eqs. (105) and (131) - (134) in the homogeneous case, which is $\Lambda=0$. We denote the homogeneous solution with a corner ”K”. When difine
@@
$\{[\frac{1}{2}-(|Q|^{2}-m^{2}|S|^{2})](Z^{K}+(Z^{K})^{*})$
$-(|Q|^{2}-m^{2}|S|^{2})(|Z^{K}|^{2}+1)\}[(\Delta_{r}^{K}(Z^{K})_{,r})_{,r}+(\Delta_{\mu}^{K}(Z^{K})_{,\mu})_{,\mu}]$
$=\{1-2(|Q|^{2}-m^{2}|S|^{2})((Z^{K})^{*}+1)\}[\Delta_{r}^{K}(Z^{K})_{,r}^{2}+\Delta_{\mu}^{K}(Z^{K})_{,\mu}^{2}],$ (150)